OpenStudy (anonymous):

Which of the following recursive formulas represents the same geometric sequence as the formula an = 3 * 3n - 1? A. an = 9 * an - 1 B. an = 3 * an - 1 C. an = 9 + an - 1 D. an = 3 + an - 1

OpenStudy (anonymous):

@ShadowLegendX @Nnesha

Nnesha (nnesha):

recursive formula is \[\huge\rm a _{n} = a _{n-1} \times r \]

OpenStudy (anonymous):

no parentheses but yes

Nnesha (nnesha):

\[3 \times 3n-1 ???\]

Nnesha (nnesha):

nope that suppose to be \[\huge a _{n} = 3 \times 3_{n-1}\] ???right ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yea thas rihgt

Nnesha (nnesha):

last one ?? oaky so what is ratio ??

OpenStudy (anonymous):

9? @Nnesha

Nnesha (nnesha):

nope look at the equation that i gave you what is r there

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

yes so now just replace r by 3